# thanks: Corresponding author

### 08 Mar 2026 v2

@@ Abstract
We study the non-leptonic decays of $\Omega_{b}\rightarrow\Omega_{c}P(V)$ with eight pseudoscalar and vector mesons using the naive factorization approach. We analyze all relevant topologies (the tree-level, color-suppressed, and penguin) of these decays and calculate the decay amplitude for each separately. We determine the decay rates, branching ratios and compare our results with those from other theoretical predictions. The results obtained may be useful for the analysis of the related data in both ongoing and future experiments.
@@ I Introduction
This study focuses on the two-body non-leptonic decay of $\Omega_{b}$, the heaviest singly heavy baryon, into an $\Omega_{c}$. While the decay channel $\Omega_{b}\rightarrow\Omega_{c}\pi$ has been experimentally observed with a significance of $3.3\sigma$ CDF:2014mon , other decay modes involving different mesons remain yet to be detected. This, highlighting the ongoing exploration and need for further experimental efforts in uncovering additional $\Omega_{b}$ decay processes. This decay mode has also been analyzed within various theoretical frameworks. Two main approaches realize approximate SU(3) flavor symmetry in QCD for two-body non-leptonic decays: The reducible method and the topological method. Topological diagrams, in particular, provide an intuitive representation of the strong-interaction dynamics embedded in weak decay processes, incorporating both perturbative and nonperturbative QCD effects. In two-body non-leptonic baryon decays governed by quark color-flavor dynamics, five distinct topological diagrams contribute via W-boson exchange, classified into reducible tree-level diagrams and irreducible W-exchange diagrams. These topologies reflect different mechanisms by which the weak interaction mediates the baryon decay processes. In tree-level diagrams, necessary condition is that two quarks are shared by the parent and daughter baryons and the W boson is emitted either externally or internally. The difference between them is that the quark produced by the weak transition goes into the final state baryon or meson. These diagrams can be factorized. However, the three different types of W-exchange diagrams are not factorable. Tree-level processes typically dominate, yet W-exchange contributions can amplify branching ratios by up to $30\%$ in decays like $\Lambda_{b}\to\Lambda_{c}$ Zhang:2022iun ; Ivanov:1997hi .
@@
This paper is organized as follows. In Sec. II, decay amplitudes are derived based on the relevant topological diagrams. In Sec.III, decay rates are calculated. Section IV presents numerical analyses and results, and Sec. V contains the conclusions. Two appendices are provided to present some details of the calculations.
@@ II The Decay Amplitude
The decay amplitude is obtained by sandwiching the effective Hamiltonian between the initial and the final hadronic states. Using factorization hypothesis and the Fierz identity, the amplitude can be expressed distinctly for the two topologoes: 1) For the external W-emission (color-allowed), the amplitude factorizes into the product of a meson decay constant and a baryon transition form factor, reflecting the color-favored topology in which the W boson directly produces the meson. 2) For the internal W-emission (color-suppressed), the amplitude involves color rearrangement via the Fierz transformation, leading to suppression due to color mismatch. This contribution factorizes similarly but is multiplied by $1/{N_{c}}$ color-suppression factor. Accordingly, the amplitudes can be written as:
@@
where $f_{P}$ and $f_{V}$ are the decay constants of the pseudoscalar and vector mesons, respectively. $m_{V}$ and $q^{\mu}$ denote the mass and four-momentum of the vector and pseudocalar mesons, respectively, $\epsilon^{\mu}$ is the polarization vector of the vector meson. These parameters quantify the overlap of the meson wave function at the origin and are crucial inputs for factorization-based amplitude calculations in heavy baryon decays.
@@ III DECAY RATES
where the center of mass frame momentum of final state particles is given by: $p_{c}=\frac{1}{2m_{A}}\lambda^{\frac{1}{2}}(m_{A}^{2},m_{B}^{2},m_{C}^{2})$ and integral is taken over the phase space. The decay rate is obtained by inserting the squared decay amplitude and the masses of the hadrons. The differential decay rate can also be written in terms of helicity amplitudes. It is important to note that choosing between the direct decay rate formula and the helicity formalism depends on how the hadronic current is expanded in terms of form factors. In the standard helicity expressions Faessler:2009xn ; Gutsche:2017wag , the baryonic current is parameterized using $\gamma^{\mu}$, $\sigma_{\mu\nu}q_{\nu}$ and $q^{\mu}$; with this choice, contracting with the $q^{\mu}$ coming from the meson matrix element remove the contribution of the form factor $F_{2}$ and $G_{2}$ in pseudoscalar decays, and similarly once the vector‑meson polarization is included, $F_{3}$ and $G_{3}$ drop out of the decay rate for vector mesons. In contrast, in the present work the baryonic current is expanded in terms of $\gamma^{\mu}$, $p^{\mu}$, and $p^{\prime\mu}$. With this parameterization, all three form factors $F_{1},F_{2}$, and $F_{3}$ (and analogously $G_{1},G_{2},G_{3}$) enter the scalar combinations that build the decay rate, so none of them drops out. For this reason, one must use the direct expression for the differential decay rate of spin 1/2 to spin 1/2 baryon with the factor $\frac{1}{2s+1}$:

### 13 Mar 2026 v3

@@ Abstract
We study the nonleptonic decays of $\Omega_{b}\rightarrow\Omega_{c}P(V)$ with eight pseudoscalar and vector mesons using the naive factorization approach. We analyze all relevant topologies (the tree-level, color-suppressed, and penguin) of these decays and calculate the decay amplitude for each separately. We determine the decay rates, branching ratios and compare our results with those from other theoretical predictions. The results obtained may be useful for the analysis of the related data in both ongoing and future experiments.
@@ I Introduction
This study focuses on the two-body non-leptonic decay of $\Omega_{b}$, the heaviest singly heavy baryon, into an $\Omega_{c}$. While the decay channel $\Omega_{b}\rightarrow\Omega_{c}\pi$ has been experimentally observed with a significance of $3.3\sigma$ CDF:2014mon , other decay modes involving different mesons remain yet to be detected. This, highlights the ongoing exploration and need for further experimental efforts in uncovering additional $\Omega_{b}$ decay processes. This decay mode has also been analyzed within various theoretical frameworks. Two main approaches realize approximate SU(3) flavor symmetry in QCD for two-body non-leptonic decays: The reducible method and the topological method. Topological diagrams, in particular, provide an intuitive representation of the strong-interaction dynamics embedded in weak decay processes, incorporating both perturbative and nonperturbative QCD effects. In two-body non-leptonic baryon decays governed by quark color-flavor dynamics, five distinct topological diagrams contribute via W-boson exchange, classified into reducible tree-level diagrams and irreducible W-exchange diagrams. These topologies reflect different mechanisms by which the weak interaction mediates the baryon decay processes. In tree-level diagrams, necessary condition is that two quarks are shared by the parent and daughter baryons and the W boson is emitted either externally or internally. The difference between them is that the quark produced by the weak transition goes into the final state baryon or meson. These diagrams can be factorized. However, the three different types of W-exchange diagrams are not factorable. Tree-level processes typically dominate, yet W-exchange contributions can amplify branching ratios by up to $30\%$ in decays like $\Lambda_{b}\to\Lambda_{c}$ Zhang:2022iun ; Ivanov:1997hi .
@@
This paper is organized as follows. In Sec. II, decay amplitudes are derived based on the relevant topological diagrams. In Sec.III, decay rates are calculated. Section IV presents numerical analyses and results, and Sec. V contains the conclusions. Two appendixes are provided to present some details of the calculations.
@@ II The Decay Amplitude
The decay amplitude is obtained by sandwiching the effective Hamiltonian between the initial and the final hadronic states. Using factorization hypothesis and the Fierz identity, the amplitude can be expressed distinctly for the two topologies: 1) For the external W-emission (color-allowed), the amplitude factorizes into the product of a meson decay constant and a baryon transition form factor, reflecting the color-favored topology in which the W boson directly produces the meson. 2) For the internal W-emission (color-suppressed), the amplitude involves color rearrangement via the Fierz transformation, leading to suppression due to color mismatch. This contribution factorizes similarly but is multiplied by $1/{N_{c}}$ color-suppression factor. Accordingly, the amplitudes can be written as:
@@
where $f_{P}$ and $f_{V}$ are the decay constants of the pseudoscalar and vector mesons, respectively. $m_{V}$ and $q^{\mu}$ denote the mass and four-momentum of the vector and pseudoscalar mesons, respectively, $\epsilon^{\mu}$ is the polarization vector of the vector meson. These parameters quantify the overlap of the meson wave function at the origin and are crucial inputs for factorization-based amplitude calculations in heavy baryon decays.
@@ III DECAY RATES
where the center of mass frame momentum of final state particles is given by: $p_{c}=\frac{1}{2m_{A}}\lambda^{\frac{1}{2}}(m_{A}^{2},m_{B}^{2},m_{C}^{2})$ and integral is taken over the phase space. The decay rate is obtained by inserting the squared decay amplitude and the masses of the hadrons. The differential decay rate can also be written in terms of helicity amplitudes. It is important to note that choosing between the direct decay rate formula and the helicity formalism depends on how the hadronic current is expanded in terms of form factors. In the standard helicity expressions Faessler:2009xn ; Gutsche:2017wag , the baryonic current is parameterized using $\gamma^{\mu}$, $\sigma_{\mu\nu}q_{\nu}$ and $q^{\mu}$; with this choice, contracting with the $q^{\mu}$ coming from the meson matrix element remove the contribution of the form factor $F_{2}$ and $G_{2}$ in pseudoscalar decays, and similarly once the vector‑meson polarization is included, $F_{3}$ and $G_{3}$ drop out of the decay rate for vector mesons. In contrast, in the present work the baryonic current is expanded in terms of $\gamma^{\mu}$, $p^{\mu}$, and $p^{\prime\mu}$. With this parameterization, all three form factors $F_{1},F_{2}$, and $F_{3}$ (and analogously $G_{1},G_{2},G_{3}$) enter the scalar combinations that build the decay rate, so none of them drop out. For this reason, one must use the direct expression for the differential decay rate of spin 1/2 to spin 1/2 baryon with the factor $\frac{1}{2s+1}$: